A state's license plates consist of three letters followed by four numerals, and 248 letter arrangements are not allowed. How many plates can the state issue?
Answer by jim_thompson5910(35256)

There are (26^3) - 248 = 17,328 letter arrangements

There are 10^4 = 10,000 numeral arrangements

-------------------------------------------------------

There are 17,328*10,000 = 173,280,000 different license plates possible.